7. You are washing the walls in your kitchen with a product to remove oils and grease before you paint. (oils and grease make it more difficult for paint to adhere) The chemical you are using is called TSP and the active ingredient is Na3PO4 (Kb = 5.9 *10 -3) The instructions have you dissolve1/4 cup (40 grams) into 1 gallon (3.785Liters). What is the pH of the wall cleaning solution
1
Expert's answer
2020-06-08T15:08:47-0400

Molar Mass of Na3PO4 = 3 × 23 + 31 + 64 = 164 gm/mol .

Concentration of Na3PO4 = (40)/ (164 × 3.785)

= 0.0644 M.

So, Concentration of PO43- = 0.0644 M

PO43- gives HPO42- and OH- after hydrolysis.


, If y is degree of hydrolysis of PO43- then Concentration of OH- is 0.0644 y M.

Concentration of H2PO4- is 0.0644 y M.

Concentration of PO43- is 0.0644 (1-y ) approx is equal to 0.0644 M.


So, Kb = (0.0644 y ) × (0.0644 y) / 0.0644 .

So, (0.0644 y)2 = Kb × 0.06444.

So,( 0.0644 y)2 = 0.0644 × 5.9 × 10-3 .

0.0644 y is the Concentration of OH- .

So, Concentration of OH- = 1.94 × 10-2 M..

So, POH = -logconcentration of OH-

So, POH = 1.712 (approx.)

So, PH = 14 - 1.712 = 12.287 (approx ).

At 250 c .